# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v5

@@ 4.3.3 Ensuring Timely Maintenance Operations
Predictability and Performance Guarantees. The key drawback of our SMD implementation is that it makes the synchronous DDRx interface less predictable. However, the practical performance overheads of this drawback are very low, as we demonstrate in Section 7. In a system without SMD, an already determined optimal DRAM command schedule executes as intended (without random delays) as the DDRx interface is synchronous. However, in an SMD system, an optimal schedule could have to be periodically recreated to remain optimal because the DRAM chips can reject activations and induce non-deterministic latency for memory requests. In this paper, we comprehensively describe only one out of many different possible implementations of SMD due to page limits. This implementation prioritizes simplicity. However, the DRAM standard can more strictly specify exactly when and for how long SMD chips are allowed to perform maintenance operations. This stricter implementation would make SMD chips as predictable as DDRx chips today. We hope that future work building on SMD will comprehensively address these issues. To make it more evident that SMD can be implemented in a way that preserves the predictability of the DDRx interface from the perspective of the memory controller (i.e., in a way that keeps the DRAM interface synchronous), we describe a (more) predictable SMD standardization.
@@ 5.1 Use Case 1: DRAM Refresh
In conventional DRAM, the MC periodically issues REF commands to initiate a DRAM refresh operation. This approach is inefficient due to three main reasons. First, transmitting REF commands over the DRAM command bus within the refresh period (e.g., 64, 32, or even depending on the refresh rate) consumes energy and increases the command bus utilization. Due to transmitting a REF over the command bus, the MC may delay a command to another rank in the same channel, which increases DRAM access latency [16]. Second, an entire bank becomes inaccessible while being refreshed although a REF command refreshes only a few rows in the bank. This incurs up to 50% loss in memory throughput [86]. Third, new mechanisms that reduce refresh overhead (e.g., by skipping refreshes on retention-strong rows) are difficult to implement. Even a simple optimization of enabling per-bank refresh in DDR4 requires changes to the DRAM interface and the MC.
@@
Leveraging SMD, we design two maintenance mechanisms for DRAM refresh: Fixed-Rate Refresh and Variable Refresh.
5.1.1 Fixed-Rate Refresh (SMD-FR)
SMD-FR refreshes DRAM rows in a fixed time interval (i.e., tREFI), similar to conventional DRAM refresh. It enables refresh-access parallelization without any burden on the MC beyond the changes required for implementing the SMD interface. As with any maintenance operation in SMD, the MC can activate a row (in any bank) in regions that are not locked while SMD-FR refreshes rows in a locked region. SMD-FR refreshes (Refresh Granularity) number of rows from a lock region and switches to refreshing another region to 1) limit the time for which a single lock region remains under maintenance and 2) prevent memory requests from waiting too long for the region to become available for access.
@@ 5.1.1 Fixed-Rate Refresh (SMD-FR)
SMD-FR operates independently in each DRAM bank and SMD-FR uses three counters for managing the refresh operations in a bank: pending refresh counter, lock region counter, and row address counter. Fig. 5 illustrates SMD-FR operation. The pending refresh counter is used to slightly delay refresh operations within the slack777DDR4 [50, 96, 45] allows the MC to postpone issuing up to REF commands in order to serve pending memory requests first. given in the DRAM standard until the region to be refreshed can be locked. The pending refresh counter is initially zero and SMD-FR increments it by one at the end of a tREFI interval . SMD-FR allows up to refresh operations to be accumulated in the pending refresh counter. As we explain in section 4.3.3, because the MC can keep a row open for a limited time, the pending refresh counter never exceeds the value , and this is how SMD ensures that the refresh operations do not lag behind more than the window of refresh operations.
@@
The lock region and row address counters indicate the next row to refresh. When pending refresh counter is greater than zero, SMD-FR attempts to lock the region indicated by the lock region counter every clock cycle until it successfully locks the region . In some cycles, SMD-FR may fail to lock the region either because the lock region contains an active row or the region is locked by another maintenance mechanism. When SMD-FR successfully locks the region, it initiates a refresh operation that refreshes number of sequentially-addressed rows in the lock region starting from the row indicated by the row address counter . The refresh operation is carried out similarly as in conventional DRAM, essentially by activating and precharging a row to restore the charge in its DRAM cells. Making larger increases the latency of a single refresh operation, which prolongs the time that a lock region remains unavailable for access. In contrast, small causes switching from one lock region to another very often, increasing interference with accesses. We empirically find to be a favorable design point.
@@
5.1.2 Variable Refresh (SMD-VR)
In conventional DRAM, all DRAM rows are uniformly refreshed with the same refresh period. However, the actual data retention times of different rows in the same DRAM chip greatly vary mainly due to manufacturing process variation and design-induced variation [86, 124, 85, 121, 78, 107]. In fact, only hundreds of “weak” rows across an entire DRAM chip require to be refreshed at the default refresh rate, and a vast majority of the rows can correctly operate when the refresh period is doubled or quadrupled [86]. Eliminating unnecessary refreshes to DRAM rows that do not contain weak cells can significantly mitigate the performance and energy consumption overhead of DRAM refresh [86].
We develop Variable Refresh (SMD-VR), a mechanism that refreshes different rows at different refresh rates depending on the retention time characteristics of the weakest cell in each row. SMD enables implementing SMD-VR with no further changes in the DRAM interface and the MC. Our SMD-VR design demonstrates the versatility of SMD in supporting different DRAM refresh mechanisms.
The high-level idea of SMD-VR is to group DRAM rows into multiple retention time bins and refresh a DRAM row based on the retention time bin that it belongs to. To achieve low design complexity, SMD-VR uses only two bins: 1) retention-weak rows that have retention time less than and 2) rows that have retention time more than . Inspired by RAIDR [86], SMD-VR stores the addresses of retention-weak rows using a per-bank Bloom Filter [10], which is a space-efficient probabilistic data structure for representing set membership. We assume retention-weak rows are already inserted into the Bloom Filters by the DRAM vendors during post-manufacturing tests.888Alternatively, SMD can be used to develop a maintenance mechanism that performs retention profiling. We believe SMD enables such profiling in a seamless way. Due to limited space, we leave the development and analysis of it to future work.
@@ 5.1.2 Variable Refresh (SMD-VR)
SMD-VR increments the pending refresh counter by one at the end of a tREFI interval . When the pending refresh counter is greater than zero, SMD-VR determines whether or not the number of rows, starting from the address indicated by the lock region and row address counters, are retention-weak rows by testing their row addresses using the bank’s Bloom Filter . SMD-VR refreshes the rows that are present in the Bloom Filter every time when it is their turn to be refreshed, as indicated by the lock region and row address counters. In contrast, SMD-VR refreshes the rows that are not present in the Bloom Filter only when the refresh cycle counter has a value that is multiple of , which is specified by Equation 1.
@@ 5.2 Use Case 2: RowHammer Protection
RowHammer is a DRAM reliability issue mainly caused due to the small DRAM cell size and the short cell-to-cell distance in modern DRAM chips. Repeatedly activating and precharging (i.e., hammering) a (aggressor) DRAM row creates a disturbance effect, which accelerates charge leakage and causes bit errors in the cells of a nearby (victim) DRAM row [158, 33, 54, 113, 114, 127, 151, 157].
The three major DRAM manufacturers equip their existing DRAM chips with in-DRAM RowHammer protection mechanisms, generally referred to as Target Row Refresh (TRR) [41, 31, 48]. At a high level, TRR protects against RowHammer by detecting an aggressor row and refreshing its victim rows. Because a conventional DRAM chip cannot initiate a refresh operation by itself, TRR refreshes victim rows by taking advantage of the slack time available in the refresh latency (i.e., tRFC), originally used to perform only periodic DRAM refresh [41, 31, 48]. This approach has two major shortcomings that limit the protection capability of TRR [41]. First, the limited slack time in tRFC puts a hard constraint on the number of TRR-induced refreshes. This restriction results in insufficient protection to victim rows, especially when DRAM becomes more RowHammer vulnerable with DRAM technology scaling. Second, TRR-induced refresh takes place only at a certain rate (i.e., at most on every REF ). Recent works [41, 31, 48, 24] demonstrate a variety of new RowHammer access patterns that circumvent the TRR protection in chips of all three major DRAM vendors, proving that the existing DRAM interface is not well suited to enable strong RowHammer protection especially as RowHammer becomes a bigger problem with DRAM technology scaling.
We use SMD to develop two maintenance mechanisms that overcome the limitations of the existing TRR mechanisms by initiating victim row refresh within the DRAM chip.
5.2.1 Probabilistic RowHammer Protection
Inspired by PARA [70], we implement an in-DRAM maintenance mechanism called Probabilistic RowHammer Protection (SMD-PRP). The high level idea is to refresh the nearby rows of an activated row with a small probability. PARA is proposed as a mechanisms in the MC, which makes it difficult to adopt since victim rows are not always known to the MC. SMD enables us to overcome this issue by implementing the PARA-inspired SMD-PRP mechanism completely within the DRAM chip. In addition, SMD-PRP avoids explicit ACT and PRE commands to be sent over the DRAM bus. Fig. 7 illustrates the operation of SMD-PRP.
@@ 5.2.1 Probabilistic RowHammer Protection
5.2.2 SMD-PRP with Aggressor Row Detection
SMD-PRP refreshes victim rows with a small probability on every row activation, even does so for a row that has been activated only a few times. This results in unnecessary victim refreshes, especially for high values that strengthen the RowHammer protection as DRAM cells become more vulnerable to RowHammer with technology scaling.
@@ 5.2.2 SMD-PRP with Aggressor Row Detection
5.2.3 Deterministic RowHammer Protection
SMD-PRP and SMD-PRP+ are probabilistic mechanisms that provide statistical security against RowHammer attacks with being the security parameter. On an extremely security-critical system, a deterministic RowHammer protection mechanisms can guarantees mitigation of RowHammer bit flips at all times.
We use SMD to implement SMD-DRP, a deterministic RowHammer protection mechanism based on the Graphene [116] mechanism, which uses the Misra-Gries algorithm [97] for detecting frequent elements in a data stream to keep track of frequently activated DRAM rows. Graphene is provably secure RowHammer protection mechanism with zero chance to miss an aggressor row activated more than a predetermined threshold. Different from Graphene, we implement SMD-DRP completely within a DRAM chip by taking advantage of SMD, whereas Graphene requires the memory controller to issue neighbor row refresh operations when necessary.
@@ 5.2.3 Deterministic RowHammer Protection
When SMD-DRP receives an ACT, it check if the activated row address () is already in CT . If so, SMD-DRP increments the corresponding counter in CT by one . If is not in CT, SMD-DRP finds the smallest counter value () in CT and compares it to the value of the spillover counter (SP) , which is initially zero. If SP is equal to , SMD-DRP replaces the row address corresponding to in CT with and increments the corresponding CT counter by one . If SP is smaller than , SMD-DRP increments SP by one . When a CT counter is incremented in step , SMD-DRP checks if the counter value is a multiple of , which is the maximum number of times a row can be activated without refreshing its neighbors. If so, SMD-DRP refreshes the neighbors of . To prevent the counters from overflowing, SMD-DRP resets the CT counters and SP on every tREFW interval.
To ensure that no row is activated more than without refreshing its neighbor rows, the number of CT counters () must be configured according to the following formula:
where is the maximum number of activations that the memory controller can perform within a tREFW interval in a single bank.
We refer the reader to [116] for more details on the Graphene mechanism and its security proof. The operation of SMD-DRP is similar to the operation of Graphene with the difference of SMD-DRP being implemented completely within a DRAM chip, which does not affect the underlying operation, and thus the security proof of Graphene applies to SMD-DRP.
@@ 5.3 Use Case 3: Memory Scrubbing
To mitigate the increasing bit error rates that are mainly caused by the continued DRAM technology scaling, DRAM vendors equip their DRAM chips with on-die Error Correction Codes (ECC) [120, 119, 109, 56]. On-die ECC is designed to correct a single bit error assuming that a failure mechanism is unlikely to incur more than one bit error in a codeword [119]. However, even when the assumption always holds, a failure mechanism can gradually incur two or more bit errors over time. A widely-used technique for preventing the accumulation of an uncorrectable number of bit errors is memory scrubbing. Memory scrubbing describes the process of periodically scanning the memory for bit errors in order to correct them before more errors occur.
@@
We propose SMD-based Memory Scrubbing (SMD-MS), which is an in-DRAM maintenance mechanism that periodically performs scrubbing on DRAM chips with on-die ECC. Using SMD, we implement SMD-MS without any further changes to the DRAM interface and the MC. Compared to conventional MC-based scrubbing, SMD-MS eliminates moving data to the MC by scrubbing within the DRAM chip, and thus reduces the performance and energy overheads of memory scrubbing.
@@
SMD-MS operation resembles the operation of SMD-FR. Similar to SMD-FR, SMD-MS maintains pending scrub counter, lock region counter, and row address counter. SMD-MS increments the pending scrub counter at fixed intervals of . When the pending scrub counter is greater than zero, SMD-MS attempts to lock the region indicated by the lock region counter. After locking the region, SMD-MS performs scrubbing operation on the row indicated by the row address counter. The scrubbing operation takes more time than a refresh operation as performing scrubbing on a row consists of three time consuming steps for each codeword in the row: 1) reading the codeword, 2) performing ECC decoding and checking for bit errors, and 3) encoding and writing back the new codeword into the row only when the decoded codeword contains a bit error. Refreshing a row takes , whereas scrubbing a row takes when no bit errors are detected.999We assume ECC decoding/encoding hardware is fully pipelined. In case of a bit error, writing a corrected codeword incurs additional latency. Therefore, SMD-MS keeps a region locked for much longer than SMD-FR. When the scrubbing operation is complete, SMD-MS releases the lock region and increments the lock region and row address counters, and decrements the pending scrub counter as in SMD-FR.

### v6

@@ 4.6 Impact to Request Scheduling
In this paper, we comprehensively describe only one out of many different possible implementations of SMD due to page limits. This implementation prioritizes simplicity. However, the DRAM standard can more strictly specify exactly when and for how long SMD chips are allowed to perform maintenance operations. This stricter implementation would make SMD chips as predictable as DDRx chips today. We hope that future work building on SMD will comprehensively address these issues. To make it more evident that SMD can be implemented in a way that preserves the predictability of the DDRx interface from the perspective of the memory controller (i.e., in a way that keeps the DRAM interface synchronous), we describe a (more) predictable SMD implementation.
@@ 5.1 Use Case 1: DRAM Refresh
In conventional DRAM, the MC periodically issues REF commands to initiate a DRAM refresh operation. This approach is inefficient due to two main reasons. First, transmitting REFs over the DRAM command bus within the refresh period (e.g., 64, 32, or even depending on the refresh rate) consumes energy and increases the command bus utilization. Due to transmitting a REF over the command bus, the MC may delay a command to another rank in the same channel, which increases DRAM access latency [16]. Second, an entire bank becomes inaccessible while being refreshed although a REF command refreshes only a few rows in the bank. This incurs up to 50% loss in memory throughput [17].
@@
Leveraging SMD, we design a new efficient maintenance mechanism for DRAM refresh: Fixed-Rate Refresh.
Fixed-Rate Refresh (SMD-FR). SMD-FR refreshes DRAM rows in a fixed time interval (i.e., tREFI), similar to conventional DRAM refresh. To limit the time that the MC waits for a region to get unlocked, SMD-FR refreshes (Refresh Granularity) number of rows from a lock region and switches to refreshing another region.
@@
SMD-FR (Fig. 4) operates independently in each DRAM bank and it uses three counters for managing the refresh operations: pending refresh counter, lock region counter, and row address counter. The pending refresh counter is initially zero and SMD-FR increments it by one at the end of a tREFI interval . SMD-FR allows up to refresh operations to be accumulated in the pending refresh counter.666DDR4 [78, 83, 85] allows the MC to postpone issuing up to REF commands in order to serve pending memory requests first. Because the MC can keep a row open for a limited time (section 4.3), the pending refresh counter never exceeds the value .
@@
The lock region and row address counters indicate the next row to refresh. When pending refresh counter is greater than zero, SMD-FR attempts to lock the region indicated by the lock region counter every clock cycle until it successfully locks the region . In some cycles, SMD-FR may fail to lock the region either because the lock region contains an active row or the region is locked by another maintenance mechanism. When SMD-FR successfully locks the region, it initiates a refresh operation that refreshes number of rows in the lock region starting from the row indicated by the row address counter . The refresh operation is conducted by activating and precharging a row to restore the charge in its DRAM cells. We empirically find to be a favorable design point.
@@
Variable Refresh (SMD-VR). In conventional DRAM, all rows are uniformly refreshed with the same refresh period. However, the actual data retention times of different rows in the same DRAM chip greatly vary mainly due to manufacturing process variation and design-induced variation [17, 15, 86, 38, 1, 18]. In fact, only hundreds of “weak” rows across an entire DRAM chip require to be refreshed at the default rate, and a vast majority of the rows can correctly operate when the refresh period is doubled or quadrupled [17]. Eliminating unnecessary refreshes to DRAM rows that do not contain weak cells can significantly mitigate the performance and energy consumption overhead of DRAM refresh [17].
We develop Variable Refresh (SMD-VR), a mechanism that refreshes different rows at different refresh rates depending on the retention time characteristics of the weakest cell in each row. Our SMD-VR design demonstrates the versatility of SMD in supporting different DRAM refresh mechanisms.
The key idea of SMD-VR is to group DRAM rows into multiple retention time bins and refresh a row based on the bin that it belongs to. To achieve low design complexity, SMD-VR uses only two bins: 1) retention-weak rows that have retention time less than and 2) rows that have retention time more than . Inspired by RAIDR [17], SMD-VR stores the addresses of retention-weak rows using a per-bank Bloom Filter [87], which is a space-efficient probabilistic data structure for representing set membership. We assume retention-weak rows are already inserted into the Bloom Filters by the DRAM vendors during post-manufacturing tests.777Alternatively, SMD can be used to develop a maintenance mechanism that performs retention profiling. We leave the development and analysis of it to future work.
@@
SMD-VR increments the pending refresh counter by one at the end of a tREFI interval . When the pending refresh counter is greater than zero, SMD-VR determines whether or not the number of rows, starting from the address indicated by the lock region and row address counters, are retention-weak rows by testing their row addresses using the bank’s Bloom Filter . SMD-VR refreshes the rows that are present in the Bloom Filter every time when it is their turn to be refreshed, as indicated by the lock region and row address counters. In contrast, SMD-VR refreshes the rows that are not present in the Bloom Filter only when the refresh cycle counter has a value that is multiple of , specified as .
@@ 5.2 Use Case 2: RowHammer Protection
Repeatedly activating and precharging (i.e., hammering) a (aggressor) DRAM row causes RowHammer errors in the cells of a nearby (victim) DRAM row [88, 89, 90, 91, 92, 93, 94, 95]. DRAM manufacturers equip their existing DRAM chips with in-DRAM RowHammer protection mechanisms, generally referred to as Target Row Refresh (TRR) [74, 75, 76]. At a high level, TRR protects against RowHammer by detecting an aggressor row and refreshing its victim rows. Because a conventional DRAM chip cannot initiate a refresh operation by itself, TRR refreshes victim rows by taking advantage of the slack time available in the refresh latency (i.e., tRFC), originally used to perform only periodic DRAM refresh [74, 75, 76]. Recent works [74, 75, 76, 96] demonstrate a variety of new RowHammer access patterns that circumvent the TRR protection in chips of all three major DRAM vendors, proving that the existing DRAM interface is not well suited to enable strong RowHammer protection especially as RowHammer becomes a bigger problem with DRAM technology scaling. We use SMD to develop a new maintenance mechanism that overcomes the limitations of the existing TRR mechanisms by initiating victim row refresh within the DRAM chip.
Probabilistic RowHammer Protection (PRP). Inspired by PARA [10], we implement an in-DRAM maintenance mechanism called Probabilistic RowHammer Protection (SMD-PRP). The high level idea is to refresh the nearby rows of an activated row with a small probability. PARA is proposed as a mechanisms in the MC, which makes it difficult to adopt since victim rows are not always known to the MC. SMD enables us to overcome this issue by implementing the PARA-inspired SMD-PRP mechanism completely within the DRAM chip. In addition, SMD-PRP avoids explicit ACT and PRE commands to be sent over the DRAM bus. We omit the implementation details of SMD-PRP due to space limitations.
Fig. 6 illustrates the operation of SMD-PRP.
@@
SMD-PRP with Aggressor Row Detection (SMD-PRP+). SMD-PRP refreshes victim rows with a small probability on every row activation, even does so for a row that has been activated only a few times. This results in unnecessary victim refreshes, especially for high values that strengthen the RowHammer protection as DRAM cells become more vulnerable to RowHammer with technology scaling.
@@
Deterministic RowHammer Protection (DRP). We use SMD to implement SMD-DRP, a deterministic RowHammer protection based on the Graphene [48] mechanism keeps track of frequently activated DRAM rows. Different from Graphene, we implement SMD-DRP completely within a DRAM chip, whereas Graphene requires the MC to issue neighbor row refresh operations.888 We refer the reader to [48] for more details on the Graphene mechanism and its security proof. The operation of SMD-DRP is similar to the operation of Graphene with the difference of SMD-DRP being implemented completely within a DRAM chip, which does not affect the underlying operation, and thus the security proof of Graphene applies to SMD-DRP.
SMD-PRP and SMD-PRP+ are probabilistic mechanisms that provide statistical security guarantees against RowHammer attacks with being the security parameter. On an extremely security-critical system, a deterministic RowHammer protection mechanisms that guarantees the mitigation of RowHammer bit flips at all times may be desired.
@@
When SMD-DRP receives an ACT, it checks if the activated row address () exists in CT . If so, SMD-DRP increments the corresponding CT counter by one . Otherwise, SMD-DRP finds the smallest counter value () in CT and compares it to the value of the spillover counter (SP) , which is initially zero. If SP is equal to , SMD-DRP replaces the row address corresponding to in CT with and increments the corresponding CT counter by one . If SP is smaller than , SMD-DRP increments SP by one . When a CT counter is incremented in step , SMD-DRP checks if the counter value is a multiple of , which is the maximum number of times a row can be activated without refreshing its neighbors. If so, SMD-DRP refreshes the neighbors of . To prevent the counters from overflowing, SMD-DRP resets the CT counters and SP on every tREFW interval.
To ensure that no row is activated more than without refreshing its neighbor rows, the number of CT counters () must be configured as follows:
where is the maximum number of activations that the MC can perform within a tREFW interval in a single bank. In our evaluations, we set to show that the overheads of SMD-DRP are small even for DRAM chips that are extremely vulnerable against RowHammer.
@@ 5.3 Use Case 3: Memory Scrubbing
To mitigate the increasing bit errors mainly caused by the continued DRAM technology scaling, DRAM vendors equip their DRAM chips with on-die Error Correction Codes (ECC) [99, 100, 101, 8]. On-die ECC is designed to correct a single bit error assuming that a failure mechanism is unlikely to incur more than one bit error in a codeword [100]. However, even when the assumption always holds, a failure mechanism can gradually incur two or more bit errors over time. A widely-used technique for preventing the accumulation of an uncorrectable number of bit errors is memory scrubbing. Memory scrubbing describes the process of periodically scanning the memory for bit errors in order to correct them before more errors occur.
@@
We propose SMD-based Memory Scrubbing (SMD-MS), which is an in-DRAM maintenance mechanism that periodically performs scrubbing on DRAM chips with on-die ECC.999To preserve ECC transparency and enable system-level policies that depend on ECC scrubbing information (e.g., detecting and replacing faulty DRAM chips), SMD-MS exposes critical ECC scrubbing information (e.g., the address of the DRAM row with the most corrected errors) to the memory controller via DRAM mode-status registers, similar to how ECC Error Check and Scrub does in DDR5 [72]. Compared to conventional MC-based scrubbing, SMD-MS eliminates moving data to the MC by scrubbing within the DRAM chip, and thus reduces the performance and energy overheads of memory scrubbing.
@@
SMD-MS operation resembles the operation of SMD-FR. Similar to SMD-FR, SMD-MS maintains pending scrub counter, lock region counter, and row address counter. SMD-MS increments the pending scrub counter at fixed intervals of . When the pending scrub counter is greater than zero, SMD-MS attempts to lock the region indicated by the lock region counter. After locking the region, SMD-MS performs scrubbing operation on the row indicated by the row address counter. The scrubbing operation takes more time than a refresh operation as performing scrubbing on a row consists of three time consuming steps for each codeword in the row: 1) reading the codeword, 2) performing ECC decoding and checking for bit errors, and 3) encoding and writing back the new codeword into the row only when the decoded codeword contains a bit error. Refreshing a row takes , whereas scrubbing a row takes when no bit errors are detected.101010We assume ECC decoding/encoding hardware is fully pipelined. In case of a bit error, writing a corrected codeword incurs latency. Therefore, SMD-MS keeps a region locked for longer than SMD-FR. When the scrubbing operation is complete, SMD-MS releases the lock region and increments the lock region and row address counters, and decrements the pending scrub counter as in SMD-FR.
Rank-Level Memory Scrubbing. Server memory modules typically use rank-level ECC [102, 103, 104]. We discuss why rank-level ECC and in-DRAM ECC scrubbing should be done separately but in a combined way in server modules. During a conventional DRAM RD operation, a DRAM chip with in-DRAM ECC (e.g., a DDR5 chip [72]) performs error correction for the in-DRAM ECC codeword without 1) writing the corrected codeword back and 2) keeping a record of the correction in mode-status registers [72]. Thus, in a rank-level ECC scrubbing operation, the memory controller cannot distinguish between an in-DRAM-ECC-corrected codeword and a codeword with no errors.
In a computing system that uses memory modules with in-DRAM ECC, relying only on the rank-level ECC scrubbing to prevent the accumulation of errors is costly. This is because the memory controller must write back every codeword that it reads from DRAM for every rank-level ECC scrubbing operation. Doing so could incur performance overheads due to the long data bus turn-around latency in DRAM (e.g., in DDR5 is approximately [105, 72]) and energy overheads due to increased data movement between the memory controller and the DRAM chips. Instead, using SMD, the computing system could use both rank-level ECC scrubbing and in-DRAM ECC scrubbing to 1) leverage the (typically) stronger error correction capability of rank-level ECC and 2) leverage the performance and energy efficiency of in-DRAM ECC while preventing the accumulation of errors.
5.4 Other Use Cases
To demonstrate that SMD is not limited to the three use cases we primarily focus on, we discuss two other use cases.
Online DRAM Error Profiling. SMD enables implementing a maintenance mechanism for online profiling of DRAM errors that may occur when operating with increased refresh period and reduced DRAM timing parameters [34, 17, 106, 107, 108, 1, 109, 110, 86, 38, 15, 111]. Such a maintenance mechanism can be used to explore how DRAM cell characteristics vary throughout the DRAM chip. To profile a DRAM region, the maintenance mechanism can use SMD to lock this region and prevent the MC from accessing it while the region is being profiled. Because the profiling operation can cause bit flips, the maintenance mechanism should temporarily buffer the original data that is stored in the profiled region. For example, for profiling errors that occur at reduced tRCD, the maintenance mechanism should first safely copy the data stored in a row to another storage space (e.g., an unused row or an SRAM buffer), before attempting to activate the row with reduced tRCD. The profiler mechanism can internally share the profiling results with other maintenance mechanisms (e.g., for skipping refreshes to rows with high retention time) and mechanisms for improving DRAM access latency and energy efficiency.
Processing in/near Memory. In the presence of an in-DRAM processing engine, SMD can help resolve access conflicts between the in-DRAM processing engine and the MC. To do so, SMD can treat the in-DRAM processing engine as a maintenance mechanism. The in-DRAM processing engine can use SMD to lock a DRAM region that it will operate on. Because SMD does not allow the MC to activate a row in a locked region, only the in-DRAM processing engine will have access to the locked region until it completes the processing and releases the region.